OpenStudy (anonymous):

Dan's business loses $9 each day. Which equation would be used to show how much money Dan's business has lost in a week? -9 × 7 -9 × (-7) 9 × 7 9 × (-7)

OpenStudy (anonymous):

Again you cannot have negative money. So just use the numbers you are given. There are 7 days in a week and he is losing 9 dollars for each one of those days. So the 3rd option would be the correct one as he has lost 9times7 dollars (63 dollars lost).

OpenStudy (anonymous):

If 11 × 19 = 209, which of the following is true? 209 ÷ (-19) = -11 -209 ÷ 11 = 19 11 × (-19) = 209 19 × 11 = 209

OpenStudy (anonymous):

this is a muliple question

OpenStudy (anonymous):

multiple

OpenStudy (anonymous):

When you divide a positive number by a negative number you get a negative number right? Since 11 X 19=209 that also means 209/19=11. So therefor dividing by a negative 19 will give you a negative number so the first one is true. The second one is not true because a negative divided by a positive is negative. The third one is also not true because the answer would have to be a negative number since you are multiplying a negative and a positive. The fourth one is true because 11 X 19 and 19 X 11 are the exact same thing. You can flip those back and forth all you want and they will still be correct. So the first one and the last one are the only ones that are true

OpenStudy (anonymous):

Bob owes Sally $12, but he pays her back $8. Which equation would be used to show how much money Bob still owes Sally? -12 + 8 = -12 - 8 = -8 + 12 = 8 - 12 =

OpenStudy (anonymous):

the original value is 12 dollars and again you cannot have negative money so the proper way to find out how much money he owes is to subtract the mney he gave her from the money he originally owed. Meaning that 12-8 is the proper way to find out how much he owes. Since there are not any that show that you must find another way to get the answer 4. Which ones of those 4 problems will equal to 4?

OpenStudy (anonymous):

The temperature Saturday is -13°, and on Sunday it is -4°. Which equation would be used to show the difference in temperature from Saturday to Sunday? -4 + 13 = -4 - (-13) = -4 - 13 4 + (-13)

OpenStudy (anonymous):

for a problem with a difference you must subtract. So for starters we know the difference between -13 and -4 is 9. So just find out which ones will come to an answer of 9

OpenStudy (anonymous):

thanks man that is it

OpenStudy (anonymous):

I apologize on the last one the answer will be the last one as the difference would be set up by this. -13-(-4) which would equal -9